a segment has one endpoint at (-5,4) and its midpoint at (-2,1). What are the coordinates of the other endpoint?
Answer by Cromlix(4381)

Midpoint formula:
(x1 + x2)/2 , (y1 + y2)/2
..............
(-5 + x2)/2 = -2
(-5 + x2) = -2 x 2
(-5 + x2) = -4
x2 = -4 + 5
x2 = 1
...............
(4 + y2)/2 = 1
(4 + y2) = 2 x 1
(4 + y2) = 2
y2 = 2 - 4
y2 = -2
.................
Coordinates of other endpoint = (1, -2)
Hope this helps. :-)